40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2122     Entry time: Mon Oct 19 23:14:32 2009
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: LSC timing issue 

I measured the noise spectrum of LSC_DARM_IN1 and OMC-LSC_DRIVE_EXC by using DTT,

while injecting the sin-wave into the OMC-LSC_DRIVE by AWG.

The attached are the results.

No significant differences appears between OMC and LSC in this measurement.

It means, in this measurement we can not figure out any timing noise which might be in LSC-clock.

In addition there are the noise floor, whose level does not change in each 3-figures. The level is about ~4*10^{-8} count/sqrt[Hz]

(The source of the noise floor is still under research.)

Attachment 1: SPE20Hz.png  59 kB  | Hide | Hide all
SPE20Hz.png
Attachment 2: SPE200Hz.png  64 kB  | Hide | Hide all
SPE200Hz.png
Attachment 3: SPE2kHz.png  64 kB  | Hide | Hide all
SPE2kHz.png
ELOG V3.1.3-